PARTIAL PHOTODISSOCIATION CROSS SECTIONS FOR O_{2} IN THE 1205 Å REGION

\underline{S. T. GIBSON} and B. R. LEWIS, Research School of Physical Sciences and Engineering The Australian National University, Canberra, ACT 0200, Australia.

Measurements of the photoabsorption cross section of O_{2} and the partial photodissociation cross section yielding O(^{1}D) have been made over the wavelength region of the Tanaka second band, (1, 0) E^{3}Σ_{u}^{−} − X^{3}Σ_{g}^{−}.^{a} Photodissociation in this wavelength region predominantly results in the atomic species O(^{1}D) + O(^{3}P), from predissociation via the B^{3}Σ_{u}^{−} electronic state.^{b} We have found a residual component of the cross section which does not yield O(^{1}D) and therefore arises from a different path to dissociation. The origin of this component is inferred from a coupled-channel Schrödinger equation (CSE) model, involving a manifold of electronic states and interactions previously used to explain experimental predissociation linewidths of the npπ_{u} ^{3}Σ_{u}^{+} Rydberg states.^{c} The CSE calculations give results consistent with the measured branching ratio into the O(^{1}D) and O(^{3}P) dissociation channels. The residual cross section is identified as arising from the (4, 0) 3pπ_{u} D^{3}Σ_{u}^{+} − X^{3}Σ_{g}^{−} band.^{d} This band is known to interfere strongly with the second band.^{e} We conclude that, whereas the npπ_{u} ^{3}Σ_{u}^{−} Rydberg states dissociate entirely into O(^{1}D) + O(^{3}P) due to a strong Rydbergvalence interaction, the 3pπ_{u} ^{3}Σ_{u}^{+} Rydberg state dissociates via two pathways, involving direct and indirect predissociation. The direct predissociation occurs from Rydberg-valence mixing of the ^{3}Σ_{u}^{+} states, with the electrostatic interaction an order of magnitude smaller than for the ^{3}Σ_{u}^{−} states, resulting in only O(^{3}P) products. Indirect predissociation arises from a spin-orbit interaction between the 3pπ_{u} Rydberg ^{3}Σ_{u}^{+} and ^{3}Σ_{u}^{−} states, yielding O(^{1}D). This interaction also provides the intensity of the (4, 0)D^{3}Σ_{u}^{+} − X^{3}Σ_{g}^{−} band.

aThe mixed Rydberg-valence E electronic state is also labelled B′ by some authors.

bL. C. Lee, T. G. Slanger, G. Black, and R. L. Sharpless, J. Chem. Phys. 67, 5602 (1977).

cS. S. Banerjee, Ph.D. Thesis, The Australian National University (1996).

dThe D electronic state also labelled β by some authors.

eB. R. Lewis, S. T. Gibson, M. Emami, and J. H. Carver, J. Quant. Spectrosc. Radiat. Transfer 40, 469 (1988).